# Channel Estimation for XL-MIMO Systems with Decentralized Baseband Processing: Integrating Local Reconstruction with Global Refinement

### v2

@@ VI-A Parameter Selection and Computational Complexity
Figure 5: NMSE of SBL-GNNs versus subarray number.
Fig. 5 illustrates the NMSE performance of the SBL-GNNs algorithm for various subarray configurations, with $P=N/2$. It is evident that the centralized algorithm significantly outperforms the SBL-GNNs algorithm. This disparity arises since the decentralized SBL-GNNs algorithm does not fully exploit the channel correlations among different subarrays. Additionally, the NMSE gap between the centralized algorithm and the SBL-GNNs algorithm widens as the number of subarrays increases. However, it is important to note that the computational complexity of the centralized algorithm is substantially higher than that of the SBL-GNNs algorithm. To illustrate this observation more intuitively, Fig. 6 presents the runtime of different algorithms. For fairness, all algorithms are executed on an i7-9700 CPU. Fig. 6 demonstrates that the runtime of the SBL-GNNs algorithm is shorter than that of traditional algorithms. This improvement is attributed to the powerful capability of GNNs in capturing the structure of channel characteristics. Specifically, the SBL-GNNs algorithm requires only a few iterations (we set $T=5$ in the simulations) to achieve satisfactory performance, whereas traditional algorithms often need hundreds of iterations to ensure convergence. Consequently, this significant reduction in the number of iterations substantially decreases the runtime of the SBL-GNNs algorithm.
Similarly, Fig. 6 provides the runtime of the overall algorithm. It can be seen that the runtime of the centralized algorithm is higher than that of all decentralized schemes due to the computational burden of high-dimensional matrix operations. In contrast, the proposed SBL-GNNs with channel refinement exhibits the shortest runtime in all considered scenarios, underscoring the computational efficiency of the proposed estimation scheme. Additionally, it is observed that the runtime of the overall SBL-GNNs algorithm with refinement is comparable when $M=8$ and $M=4$. Therefore, considering both performance and computational complexity, setting the number of subarrays to 4 is a practical choice for subsequent simulations.
Figure 6: Computational complexity evaluation. (a) The algorithm runtime in LPUs. (b) The overall algorithm runtime.
@@ VI-B Performance Evaluation of the SBL-GNNs Algorithm
In this subsection, we investigate the NMSE performance of different sparse recovery algorithms executed on the LPUs. Fig. 7 examines the NMSE performance of various algorithms versus SNR with $M=4$, $N=32$, and $P=16$. It is evident that the StdSBL algorithm exhibits the worst estimation performance across all considered SNR ranges. This disparity arises since the StdSBL algorithm employs a two-layer sparse prior, assuming independence among different angular and delay domain coefficients, which prevents the full exploitation of correlations or clustered sparsity among the coefficients. In contrast, algorithms based on clustered-sparse priors significantly outperform the StdSBL algorithm.
@@
Figure 7: Performance evaluation of the proposed SBL-GNNs algorithm. (a) NMSE versus SNR; (b) NMSE versus pilot symbols.
@@
Furthermore, the performance comparison of the proposed SBL-GNNs algorithm and the benchmarks at different pilot symbols $P$ is investigated in Fig. 7 with $SNR=5$dB. The number of pilot symbols $P$ increases from 12 to 32, so the compressive ratio $PN_{\mathrm{RF}}/N$ correspondingly increases from 0.375 to 1. The estimation performance of all algorithms improves with the increase in the number of pilot symbols. Notably, the proposed SBL-GNNs algorithm consistently demonstrates superior estimation performance across all considered lengths of pilot symbols. Therefore, the SBL-GNNs algorithm provides a low pilot overhead estimation scheme while improving estimation performance.
@@ VI-C Performance Evaluation of the Refinement Algorithm
Figure 8: Performance evaluation for overall algorithms. (a) NMSE gap versus SNR; (b) NMSE versus SNR; (c) NMSE versus pilot symbols.
@@
To further evaluate the performance of the overall algorithm, Fig. 8 plots the NMSE performance of the overall algorithm versus SNR with $P=16$. It can be seen that, owing to the centralized processing, the centralized algorithms achieve the best estimation performance, serving as a lower bound. Meanwhile, the performance of the proposed decentralized SBL-GNNs with refinement is closest to that of the centralized algorithm, with a gap of about 1 dB. In contrast, the performance of the best algorithm in the benchmark is approximately 2 dB less than that of the centralized algorithm. A similar superiority is observed in Fig. 8, where, when the number of pilot symbols exceeds 25, the performance of the proposed algorithm even surpasses that of the centralized algorithm. Therefore, the proposed estimation algorithm provides a low-overhead estimation scheme while ensuring robust estimation performance.
@@ VII Conclusion
According to the equivalent AWGN model in (24), $p(\mathbf{x}_{m}^{(t)},\boldsymbol{\gamma}_{m}^{(t)},\mathbf{s}_{m}^{(t)}|\boldsymbol{\mu}_{m}^{(t)})$ can be given by
$p\left(\mathbf{x}_{m}^{(t)},\boldsymbol{\gamma}_{m}^{(t)},\mathbf{s}_{m}^{(t)}|\boldsymbol{\mu}_{m}^{(t)}\right)$ (A.1)
$\propto$ $p\left(\boldsymbol{\mu}_{m}^{(t)}|\mathbf{x}_{m}^{(t)}\right)p\left(\mathbf{x}_{m}^{(t)};\boldsymbol{\gamma}_{m}^{(t)},\mathbf{s}_{m}^{(t)}\right),$
where $p(\boldsymbol{\mu}_{m}^{(t)}|\mathbf{x}_{m}^{(t)})$ and $p(\mathbf{x}_{m}^{(t)},\boldsymbol{\gamma}_{m}^{(t)},\mathbf{s}_{m}^{(t)})$ denote the likelihood and prior distributions, respectively. In the following, we first factorize the likelihood and prior distributions. Then, we further consider the factorization of the posterior distribution. In the terms of $p(\boldsymbol{\mu}_{m}^{(t)}|\mathbf{x}_{m}^{(t)})$, we have (A.2), where $\psi_{\mathrm{like}}(x_{m,n}^{(t)})\triangleq\exp\{2\mathrm{Re}((\boldsymbol{\lambda}_{m}^{(t)})^{\mathrm{H}}\mathbf{v}_{m,n}^{(t)}{x}_{m,n}^{(t)})-v^{n,n}_{m}(t)\lvert x_{m,n}^{(t)}\rvert^{2}\}$ denotes the self-potential of $x_{m,n}^{(t)}$; $\phi_{\mathrm{like}}(x_{m,n}^{(t)},x_{m,k}^{(t)})\triangleq\exp\{-v^{n,k}_{m}(t)x_{m,n}^{(t)}(x_{m,k}^{(t)})^{*}\}$ denotes the pair-potential between $x_{m,n}^{(t)}$ and $x_{m,k}^{(t)}$; $\mathbf{v}_{m,n}^{(t)}$ and $v^{n,n}_{m}(t)$ indicate the $n$-th and $(n,k)$-th entry of $\mathbf{V}_{m}^{(t)}=(\boldsymbol{\Sigma}_{m}^{(t)})^{-1}$, respectively.
$p\left(\boldsymbol{\mu}_{m}^{(t)}|\mathbf{x}_{m}^{(t)}\right)$ $\propto\exp\left\{-\left(\mathbf{x}_{m}^{(t)}-\boldsymbol{\lambda}_{m}^{(t)}\right)^{\mathrm{H}}\left(\boldsymbol{\Sigma}_{m}^{(t)}\right)^{-1}\left(\mathbf{x}_{m}^{(t)}-\boldsymbol{\mu}_{m}^{(t)}\right)\right\}$ (A.2)
$=\prod_{n=1}^{N}\exp\left\{2\mathrm{Re}\left((\boldsymbol{\mu}_{m}^{(t)})^{\mathrm{H}}\mathbf{v}_{m,n}^{(t)}{x}_{m,n}^{(t)}\right)-v^{n,n}_{m}(t)\lvert x_{m,n}^{(t)}\rvert^{2}\right\}\prod_{n=1}^{N}\prod_{k\neq n}^{N}\exp\left\{-v^{n,k}_{m}(t)x_{m,n}^{(t)}\left(x_{m,k}^{(t)}\right)^{*}\right\}$
$=\prod_{n=1}^{N}\psi_{\mathrm{like}}\left(x_{m,n}^{(t)}\right)\prod_{n=1}^{N}\prod_{k\neq n}^{N}\phi_{\mathrm{like}}\left(x_{m,n}^{(t)},x_{m,k}^{(t)}\right),$
$p\left(\mathbf{x}_{m}^{(t)},\boldsymbol{\gamma}_{m}^{(t)},\mathbf{s}_{m}^{(t)}\right)$ $=\prod_{n=1}^{N}p\left(x_{m,n}^{(t)}|\gamma_{m,n}^{(t)}\right)p\left(\gamma_{m,i}^{(t)}|s_{m,n}^{(t)}\right)\psi\left(s_{m,n}^{(t)}\right)\prod_{n=1}^{N}\prod_{k\neq n}^{N}\phi\left(s_{m,n}^{(t)},s_{m,k}^{(t)}\right)$ (A.3)
$=\prod_{n=1}^{N}\psi_{\mathrm{pri}}\left(\alpha_{m},x_{m,n}^{(t)}\right)\prod_{n=1}^{N}\prod_{k\neq n}^{N}\phi_{\mathrm{pri}}\left(\beta_{m},x_{m,n}^{(t)},x_{m,k}^{(t)}\right),$
As for $p(\mathbf{x}_{m}^{(t)})$, according to (25)-(28), we have (A.3), where $\psi_{\mathrm{pri}}(\alpha_{m},x_{m,n}^{(t)})$ and $\phi_{\mathrm{pri}}(\beta_{m},x_{m,n}^{(t)},x_{m,k}^{(t)})$ implicitly denote pair-potential and self-potential functions for prior probability distribution $p(\mathbf{x}_{m}^{(t)})$ with $\alpha_{m}$ and $\beta_{m}$ indicating the parameters of the MRF. Combining (A.2) and (A.3), (30) can be obtained.

### v4

@@ VI-A Parameter Selection and Computational Complexity
Fig. 5 illustrates the NMSE performance of the SBL-GNNs algorithm for various subarray configurations, with $P=N/2$. It is evident that the centralized algorithm significantly outperforms the SBL-GNNs algorithm. This disparity arises since the decentralized SBL-GNNs algorithm does not fully exploit the channel correlations among different subarrays. Additionally, the NMSE gap between the centralized algorithm and the SBL-GNNs algorithm widens as the number of subarrays increases. However, it is important to note that the computational complexity of the centralized algorithm is substantially higher than that of the SBL-GNNs algorithm. To illustrate this observation more intuitively, Fig. 6 presents the runtime of different algorithms. For fairness, all algorithms are executed on an i7-9700 CPU. Fig. 6(a) demonstrates that the runtime of the SBL-GNNs algorithm is shorter than that of traditional algorithms. This improvement is attributed to the powerful capability of GNNs in capturing the structure of channel characteristics. Specifically, the SBL-GNNs algorithm requires only a few iterations (we set $T=5$ in the simulations) to achieve satisfactory performance, whereas traditional algorithms often need hundreds of iterations to ensure convergence. Consequently, this significant reduction in the number of iterations substantially decreases the runtime of the SBL-GNNs algorithm.
(a) Algorithm runtime in LPUs.
(b) Overall algorithm runtime
Figure 6: Computational complexity evaluation.
Similarly, Fig. 6(b) provides the runtime of the overall algorithm. It can be seen that the runtime of the centralized algorithm is higher than that of all decentralized schemes due to the computational burden of high-dimensional matrix operations. In contrast, the proposed SBL-GNNs with channel refinement exhibits the shortest runtime in all considered scenarios, underscoring the computational efficiency of the proposed estimation scheme. Additionally, it is observed that the runtime of the overall SBL-GNNs algorithm with refinement is comparable when $M=8$ and $M=4$. This is because the total runtime is determined by the combined execution time of Algorithm 1 and Algorithm 2, under the assumption of wired fronthaul transmission and perfect synchronization. Notably, since Algorithm 2 exhibits a significantly higher computational cost compared to Algorithm 1, the overall runtime is predominantly influenced by the Bayesian refinement process. As a result, variations in $M$ have a limited impact on the total runtime. Considering both estimation performance and computational complexity, we adopt $M=4$ as a practical choice for subsequent simulations.
@@ VI-B Performance Evaluation of the SBL-GNNs Algorithm
(a) NMSE versus SNR.
(b) NMSE versus pilot symbols.
Figure 7: Performance evaluation of the SBL-GNNs algorithm.
In this subsection, we investigate the NMSE performance of different sparse recovery algorithms executed on the LPUs. Fig. 7(a) examines the NMSE performance of various algorithms versus SNR with $M=4$, $N=32$, and $P=16$. It is evident that the StdSBL algorithm exhibits the worst estimation performance across all considered SNR ranges. This disparity arises since the StdSBL algorithm employs a two-layer sparse prior, assuming independence among different angular and delay domain coefficients, which prevents the full exploitation of correlations or clustered sparsity among the coefficients. In contrast, algorithms based on clustered-sparse priors significantly outperform the StdSBL algorithm.
@@
Furthermore, the performance comparison of the proposed SBL-GNNs algorithm and the benchmarks at different pilot symbols $P$ is investigated in Fig. 7(b) with $SNR=5$dB. The number of pilot symbols $P$ increases from 12 to 32, so the compressive ratio $PN_{\mathrm{RF}}/N$ correspondingly increases from 0.375 to 1. The estimation performance of all algorithms improves with the increase in the number of pilot symbols. Notably, the proposed SBL-GNNs algorithm consistently demonstrates superior estimation performance across all considered lengths of pilot symbols. Therefore, the SBL-GNNs algorithm provides a low pilot overhead estimation scheme while improving estimation performance.
@@ VI-C Performance Evaluation of the Refinement Algorithm
Figure 8: NMSE gap versus SNR.
@@
(a) NMSE versus SNR.
(b) NMSE versus pilot symbols.
Figure 9: Performance evaluation for overall algorithms.
To further evaluate the performance of the overall algorithm, Fig. 9(a) plots the NMSE performance of the overall algorithm versus SNR with $P=16$. It can be seen that, owing to the centralized processing, the centralized algorithms achieve the best estimation performance, serving as a lower bound. Meanwhile, the performance of the proposed decentralized SBL-GNNs with refinement is closest to that of the centralized algorithm, with a gap of about 1 dB. In contrast, the performance of the best algorithm in the benchmark is approximately 2 dB less than that of the centralized algorithm. A similar superiority is observed in Fig. 9(b), where, when the number of pilot symbols exceeds 25, the performance of the proposed algorithm even surpasses that of the centralized algorithm. Therefore, the proposed estimation algorithm provides a low-overhead estimation scheme while ensuring robust estimation performance.
$p(\boldsymbol{\mu}_{m}^{(t)}|\mathbf{x}_{m})$ $\propto\exp\left\{-\left(\mathbf{x}_{m}-\boldsymbol{\mu}_{m}^{(t)}\right)^{\mathrm{H}}\left(\boldsymbol{\Sigma}_{m}^{(t)}\right)^{-1}\left(\mathbf{x}_{m}-\boldsymbol{\mu}_{m}^{(t)}\right)\right\}$ (A.2)
$=\prod_{n\in\mathcal{V}}\exp\left\{2\mathrm{Re}\left((\boldsymbol{\mu}_{m}^{(t)})^{\mathrm{H}}\mathbf{v}_{m,n}^{(t)}{x}_{m,n}\right)-v^{n,n}_{m}(t)\lvert x_{m,n}\rvert^{2}\right\}\prod_{n\in\mathcal{V}}\prod_{k\neq n}\exp\left\{-v^{n,k}_{m}(t)x_{m,n}x_{m,k}^{*}\right\}$
$=\prod_{n\in\mathcal{V}}\psi_{\mathrm{like}}\left(x_{m,n}\right)\prod_{n\in\mathcal{V}}\prod_{k\neq n}\phi_{\mathrm{like}}\left(x_{m,n},x_{m,k}\right),$
$p\left(\mathbf{x}_{m};\boldsymbol{\gamma}_{m},\mathbf{s}_{m}\right)$ $=\prod_{n\in\mathcal{V}}p\left(x_{m,n}|\gamma_{m,n}\right)p\left(\gamma_{m,n}|s_{m,n}\right)\psi\left(s_{m,n}\right)\prod_{n\in\mathcal{V}}\prod_{k\neq n}\phi\left(s_{m,n},s_{m,k}\right)$ (A.3)
$=\prod_{n\in\mathcal{V}}\psi_{\mathrm{pri}}\left(\alpha_{m},x_{m,n}\right)\prod_{n\in\mathcal{V}}\prod_{k\neq n}\phi_{\mathrm{pri}}\left(\beta_{m},x_{m,n},x_{m,k}\right),$
@@ VII Conclusion
According to the equivalent AWGN model in (24), $p(\mathbf{x}_{m},\boldsymbol{\gamma}_{m},\mathbf{s}_{m}|\boldsymbol{\mu}_{m}^{(t)})$ can be given by
$p(\mathbf{x}_{m},\boldsymbol{\gamma}_{m},\mathbf{s}_{m}|\boldsymbol{\mu}_{m}^{(t)})\propto p(\boldsymbol{\mu}_{m}^{(t)}|\mathbf{x}_{m})p\left(\mathbf{x}_{m};\boldsymbol{\gamma}_{m},\mathbf{s}_{m}\right),$ (A.4)
where $p(\boldsymbol{\mu}_{m}^{(t)}|\mathbf{x}_{m})$ and $p\left(\mathbf{x}_{m};\boldsymbol{\gamma}_{m},\mathbf{s}_{m}\right)$ denote the likelihood and prior distributions, respectively. In the following, we first factorize the likelihood and prior distributions. Then, we further consider the factorization of the posterior distribution. In the terms of $p(\boldsymbol{\mu}_{m}^{(t)}|\mathbf{x}_{m})$, we have (A.2), where $\psi_{\mathrm{like}}(x_{m,n})\triangleq\exp\{2\mathrm{Re}((\boldsymbol{\lambda}_{m}^{(t)})^{\mathrm{H}}\mathbf{v}_{m,n}^{(t)}{x}_{m,n})-v^{n,n}_{m}(t)\lvert x_{m,n}\rvert^{2}\}$ denotes the self-potential of $x_{m,n}$; $\phi_{\mathrm{like}}(x_{m,n},x_{m,k})\triangleq\exp\{-v^{n,k}_{m}(t)x_{m,n}x_{m,k}^{*}\}$ denotes the pair-potential between $x_{m,n}$ and $x_{m,k}$; $\mathbf{v}_{m,n}^{(t)}$ and $v^{n,k}_{m}(t)$ indicate the $n$-th column and $(n,k)$-th entry of $\mathbf{V}_{m}^{(t)}=(\boldsymbol{\Sigma}_{m}^{(t)})^{-1}$, respectively.
As for $p(\mathbf{x}_{m};\boldsymbol{\gamma}_{m},\mathbf{s}_{m})$, according to (25)-(28), we have (A.3), where $\psi_{\mathrm{pri}}(\alpha_{m},x_{m,n})$ and $\phi_{\mathrm{pri}}(\beta_{m},x_{m,n},x_{m,k})$ implicitly denote pair-potential and self-potential functions for the prior probability distribution $p(\mathbf{x}_{m};\boldsymbol{\gamma}_{m},\mathbf{s}_{m})$ with $\alpha_{m}$ and $\beta_{m}$ indicating the parameters of the MRF. Combining (A.2) and (A.3), (29) can be obtained.